# Gas-phase metallicity gradients in galaxies at Key Words.: Galaxies: high-redshift ā Galaxies: abundances ā Galaxies: ISM ā Galaxies: evolution ā Techniques: imaging spectroscopy ā Techniques: high angular resolution1 Introduction

### v1

@@ 7 Gas-phase metallicity gradients
Source name | [dex kpc] |
A2744-YD4 | 0.14 |
A2744-YD1 | ā0.05 |
BDF-3299 | ā0.03 |
COSMOS24108 | 0.03 |
To more robustly constrain the spatially resolved metallicity, we extracted the line fluxes from spatially integrated spectra from concentric radial annuli, as described in Sect. 2.2. This allowed us to increase the S/N of the emission lines at large radii and determine the dust extinction from the faint Balmer lines. We thus determined the gas-phase metallicity as a function of the distance from the centre of each source, defined as explained in the following.
The resulting metallicity gradients are reported in Fig. 6. In the system A2744-YD4, we determined the metallicity gradients only for the two brightest sources YD4 and YD1 (centred around their continuum peak) since the metallicity measurements for the other sources suffered from large uncertainties due to low S/N of emission lines. For BDF-3299, since we could have only extracted two radial points for each spatial component (BDF-3299-a, b, and c) due to the compactness of this system, we obtained the metallicity gradient profile for the whole BDF-3299 system, centred on the brightest source in line emission, BDF-3299-b. For the COSMOS24108 system, given that the line emission peaks between the two main continuum components, COSMOS24108-a and COSMOS24108-b, and not on either of them (see Fig. 4), we opted to extract a single set of concentric annular apertures centred on the line emission peak, rather than two sets centred on each of the continuum sources (since we considered this latter approach not very sensible in this case). We also report the metallicity gradient from two additional smaller apertures centred on the emission line clumps to the SE of the main system, COSMOS24108-Ea and COSMOS24108-Eb.
@@ 8 Discussion
We found flat gradients within the uncertainties for the sources in A2744-YD4, BDF-3299, and COSMOS24108 analysed in this work. As introduced in Sect. 1 and sketched in Fig.Ā 1, negative (radially decreasing) gradients are expected as a consequence of the inside-out galaxy formation scenario, while inverted (radially increasing) gradients may arise as a consequence of pristine gas accretion towards the central regions. Possible explanations for flat radial gradients are radial gas mixing processes, occurring as a consequence of mergers, SN-driven winds and/or large-scale gas circulation (āgalactic fountainsā, i.e. metal-loaded outflows expelled from the galaxy and re-accreted in the external regions), which redistribute the gas in the galaxy and wash out any pre-existing radial gradient of metallicity.
All the targets in this work show multiple sources within a few kpc and disturbed morphologies. This indicates that they are experiencing, or have experienced, interactions and galaxy merging processes, which are expected to be more frequent in the early Universe according to cosmological simulations (see e.g. Kohandel etĀ al., 2020; Pallottini & Ferrara, 2023). Specifically, galaxies at 6 with stellar masses 510 living in dense environments, similar to those studied in this work (see TableĀ 3), are expected to have already experienced multiple merger events (Gelli etĀ al., 2020). On the other hand, while in the low-spectral resolution ( 100) NIRSpec PRISM data analysed in this work the emission lines are spectrally unresolved, [Oāiii] and H do not show evidence for asymmetric wings or any complex line profile indicative of outflows in the higher resolution ( 2700) NIRSpec IFU or MSA data from Hashimoto etĀ al. (2023) and Morishita etĀ al. (2023), respectively, for the case of YD4, YD1, and s1 in the A2744-YD4 system (YD4 having both IFU and MSA 2700 data). Therefore, this indicates that strong outflows are not occurring at present in this system. However, in principle we cannot exclude that the redistribution of metals across the galaxy may also be the result of past intense SN winds mixing the ISM or re-accretion in the outer regions of metal-enriched gas expelled by past galactic outflows, whose trace is absent in present-day spectra. All in all, mergers and/or possibly either past galactic outflows or SN wind mixing seem the most likely mechanism driving the observed flat metallicity gradients.
@@ 8.1 Metallicity gradients across cosmic time
In Fig. 7, we report the metallicity gradients from a compilation of observational studies, from = 0 up to 8 (this study), namely the highest redshift probed so far. The flat gradients within the uncertainties that we find at 6ā8 are compatible with the other currently available measured gradients at these redshifts, namely Vallini etĀ al. (2024) from FIR lines with ALMA and Arribas etĀ al. (2023) from rest-frame optical lines with JWST/NIRSpec. Specifically, the results from Vallini etĀ al. (2024) showcase the potential for synergies between spatially resolved observations with JWST and ALMA, though the spatial resolutions may be quite different ( 0.2ā0.4 kpc in our case versus 1.5 kpc in the case of Vallini etĀ al. 2024).
We also report the predictions for the evolution of metallicity gradients with redshift from cosmological simulations. Some of them predict steeper negative gradients with increasing redshift (MUGS, with normal feedback, Gibson etĀ al. 2013; TNG50, Hemler etĀ al. 2021), while others predict median flatter gradients at higher though with increasing scatter (MAGICC, with enhanced feedback, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022).
Unfortunately, predictions for the cosmic evolution of metallicity gradients from simulations are limited to 3 so far. Nevertheless, if we assume that the extrapolation at higher follows the same trend which is in place up to 3, we can compare the predictions with our measurements at 6ā8. Specifically, a rough linear extrapolation of the trends from TNG50 (Hemler etĀ al., 2021) and MUGS (Gibson etĀ al., 2013) would give a metallicity gradient at 8 of ā0.2 dexĀ kpc and ā0.5 dexĀ kpc, respectively, which are both incompatible with our estimates considering the uncertainties (Fig. 7).
Our results, showing flat gradients within the uncertainties, are more compatible with expectations from those cosmological simulations predicting flatter gradients at high (MaGICC, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022). This conclusion is based on the extrapolation of current model predictions, therefore we stress the need for updated metallicity gradient predictions from simulations at 3 in order to match the redshifts reached by current observations.
@@
In summary, the flat radial gradients and the asymmetric spatial distribution of metallicity observed in some cases seem to support the merging scenario, though ISM mixing due to past SN winds or re-accretion in the outer regions of metal-enriched gas from past galactic outflows (galactic fountains) are also a viable mechanism to explain the flatness of the gradients. Spatially asymmetric low-metallicity accretion from the CGM and IGM could also contribute to produce non-azimuthally symmetric metallicities as suggested by Arribas etĀ al. (2023).
@@ 9 Conclusions
The targets have very disturbed morphologies in both the stellar continuum and ionised gas line emission. We identify several spatial components concentrated within a few kiloparsecs in all the three systems. We obtained the main integrated properties of each spatial component in each target, specifically stellar mass from SED fitting, SFR from H or H, and gas-phase metallicity by means of flux ratios of rest-frame optical emission lines which we modelled with Gaussian functions. We found log() 7.6ā9.3, SFRs 1ā15 Ā yr, and gas-phase metallicities 12+log(O/H) 7.7ā8.3 by extracting integrated spectra from circular apertures centred on each spatial component.
We compared the stellar masses and gas-phase metallicities of the targets with the mass-metallicity relations (MZRs) inferred observationally at different redshifts and with those predicted by cosmological simulations. In general, the sources in the systems studied in this work are consistent with the distribution of the highest- galaxies to date ( 3ā10) from JADES, CEERS, and EROs samples in the mass-metallicity plane. Nevertheless, most of the sources in A2744-YD4 and COSMOS24108 lie at the upper end of this 3ā10 distribution, in terms of metallicity at a given stellar mass, and are closer to the best-fit MZRs measured at 2ā3 (e.g. MOSDEF and GLASS surveys) than to the MZR at 3ā10. Relative to the MZRs from cosmological simulations, the sources in the three systems studied in this work are in good agreement with the slopes predicted at 6ā8 by most simulations.
@@
Cosmological simulations make very different predictions regarding the cosmic evolution of metallicity gradients, though they are limited to 2ā3. Some of them predict steeper negative gradients with increasing redshift (e.g. MUGS and TNG50), while others predict median flatter gradients at higher (e.g. MaGICC, FIRE, and EAGLE). The results of this work, indicating mainly flat gradients, are more compatible with the predictions from this latter group of simulations, when considering their extrapolation at 6ā8.
All in all, the results of this work in terms of the MZR and gas-phase metallicity gradients at 6ā8 provide important constraints to guide future cosmological simulations. In particular, they urge for specific predictions on the cosmic evolution of metallicity gradients and metallicity maps out to redshift of 8, given that such predictions are currently limited to 3.

### v2

@@ 7 Gas-phase metallicity gradients
| Source name | [dex kpc-1] | [dex ] |
| A2744-YD4 | 0.14 | 0.10 |
| A2744-YD1 | ā0.05 | ā0.04 |
| BDF-3299 | ā0.03 | ā0.02 |
| COSMOS24108 | ā0.03 | ā0.04 |
To more robustly constrain the spatially resolved metallicity, we extracted the line fluxes from spatially integrated spectra from concentric radial annuli, as described in Sect. 2.2. This allowed us to increase the S/N of the emission lines at large radii and determine the dust extinction from the faint Balmer lines. We thus determined the gas-phase metallicity as a function of the distance from the centre of each source, as defined in the following.
The resulting metallicity gradients are reported in Fig. 6. In the system A2744-YD4, we determined the metallicity gradients only for the two brightest sources YD4 and YD1 (centred around their continuum peak) since the metallicity measurements for the other sources suffered from large uncertainties due to low S/N of emission lines. For BDF-3299, since we could have only extracted two radial points for each spatial component (BDF-3299-a, b, and c) due to the compactness of this system, we obtained the metallicity gradient profile for the whole BDF-3299 system, centred on the brightest source in line emission, BDF-3299-b. For the COSMOS24108 system, we centred the gradient around the continuum emission peak, corresponding to the spatial component COSMOS24108-a (Fig.Ā 4, left panel). We masked the region surrounding the low-metallicity [Oāiii] clump COSMOS24108-Ea (shaded circle in Fig.Ā 6), since it has a markedly low metallicity (see Fig.Ā 4, right panel) and can be clearly spatially separated from the rest of the system. Given that the line emission peaks between the two main continuum components, COSMOS24108-a and COSMOS24108-b, and not on either of them (see Fig. 4, central panel), we also extracted the gradient centred around the [Oāiii] peak, instead of around the continuum. This is reported in Fig.Ā C in the Appendix666The Appendix is available at https://doi.org/10.5281/zenodo.13327942.. The two gradients are not statistically different within the uncertainties. In Fig.Ā C we also report the metallicity gradient from two additional smaller apertures centred on the emission line clumps to the SE of the main system, COSMOS24108-Ea and COSMOS24108-Eb.
@@
We also estimate the gradients relative to the effective (half-light) radius, , as sometimes done in observational (e.g. Belfiore etĀ al. 2017) and theoretical studies (e.g. MollĆ” etĀ al. 2019); in this case, the gradient should not depend on other properties of galaxies (as suggested by e.g. Garnett etĀ al. 1997). We estimated the effective radius with a 2D Gaussian modelling of the continuum emission for the whole BDF-3299 and COSMOS24108 systems (including all the spatial components in continuum), and for the single YD4 and YD1 components, for consistency with the way metallicity gradients were calculated; for the same reason, we employed the wavelength-dependent spatially smoothed data cube, as done for the gradients. We stress that, therefore, the thus obtained should not be considered as the intrinsic ones (i.e. de-convolved for the spatial resolution of the observations), and, in the case of BDF-3299 and COSMOS24108, they are not the of the single spatial clumps. The obtained are 0.74, 0.70, 0.75, and 1.2 kpc for YD4, YD1, BDF-3299, and COSMOS24108, respectively. The resulting gradients are reported in TableĀ 4. These are flat, and in any case ā0.1Ā dexĀ when considering the uncertainties; in the local Universe, such gradients are typical of low-mass galaxies ( 9.5ā10; Belfiore etĀ al. 2017).
@@ 8 Discussion
8.1 Driving mechanisms of the observed gradients
We find flat gradients within the uncertainties for the sources in A2744-YD4, BDF-3299, and COSMOS24108 analysed in this work. As introduced in Sect. 1 and sketched in Fig.Ā 1, negative (radially decreasing) gradients are expected as a consequence of the inside-out galaxy formation scenario, while inverted (radially increasing) gradients may arise as a consequence of pristine gas accretion towards the central regions. Possible explanations for flat radial gradients are radial gas mixing processes, occurring as a consequence of mergers, SN-driven winds and/or large-scale gas circulation (galactic fountains, i.e. metal-loaded outflows expelled from the galaxy and re-accreted in the external regions), which redistribute the gas in the galaxy and wash out any pre-existing radial gradient of metallicity.
All the targets in this work show multiple sources within a few kpc and disturbed morphologies. This indicates that they are experiencing, or have experienced, interactions and galaxy merging processes, which are expected to be more frequent in the early Universe according to cosmological simulations (see e.g. Kohandel etĀ al., 2020; Pallottini & Ferrara, 2023). Specifically, galaxies at 6 with stellar masses 5108 living in dense environments, similar to those studied in this work (see TableĀ 2), are expected to have already experienced multiple merger events (Gelli etĀ al., 2020). On the other hand, while in the low-spectral resolution ( 100) NIRSpec PRISM data analysed in this work the emission lines are spectrally unresolved, [Oāiii] and H do not show evidence for asymmetric wings or any complex line profile indicative of outflows in the higher resolution ( 2700) NIRSpec IFU or MSA data from Hashimoto etĀ al. (2023) and Morishita etĀ al. (2023), respectively, for the case of YD4, YD1, and s1 in the A2744-YD4 system (YD4 having both IFU and MSA 2700 data). Therefore, this indicates that strong outflows are not occurring at present in this system. However, in principle we cannot exclude that the redistribution of metals across the galaxy may also be the result of past intense SN winds mixing the ISM or re-accretion in the outer regions of metal-enriched gas expelled by past galactic outflows, whose trace is absent in present-day spectra. All in all, mergers and/or possibly either past galactic outflows or SN wind mixing seem the most likely mechanism driving the observed flat metallicity gradients.
@@ 8.2 Metallicity gradients across cosmic time
In Fig. 7, we report the metallicity gradients from a compilation of observational studies, from = 0 up to 8 (this study), namely the highest redshift probed so far. We report the gradients in dexĀ kpc-1, given that most values reported in literature are in these units rather then in dexĀ . The flat gradients within the uncertainties that we find at 6ā8 are compatible with the other currently available measured gradients at these redshifts, namely Vallini etĀ al. (2024) from FIR lines with ALMA and Arribas etĀ al. (2023) from rest-frame optical lines with JWST/NIRSpec. Specifically, the results from Vallini etĀ al. (2024) showcase the potential for synergies between spatially resolved observations with JWST and ALMA, though the spatial resolutions may be quite different ( 0.2ā0.4 kpc in our case versus 1.5 kpc in the case of Vallini etĀ al. 2024).
We also report the predictions for the evolution of metallicity gradients with redshift from cosmological simulations. Some of them predict steeper negative gradients with increasing redshift (MUGS, with normal feedback, Gibson etĀ al. 2013; TNG50, Hemler etĀ al. 2021), in some cases extremely steep (Taylor & Kobayashi, 2017, for which the case of a = 12 galaxy at the center of a cluster with no AGN is shown), while others predict median flatter gradients at higher though with increasing scatter (MAGICC, with enhanced feedback, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022). Predictions from chemical evolution models, also displayed in Fig.Ā 7, are very heterogeneous as well, with little to no evolution with redshift (Mott etĀ al. 2013 for the Milky Way, with variable SFE, and Sharda etĀ al. 2021, for which the = 11.1 case is shown) and slightly to steeply increasing gradients with (MollĆ” etĀ al. 2019 and Kubryk etĀ al. 2015, respectively, both for the Milky Way), as well as gradients inverting at z 1 (Mott etĀ al. 2013 with constant SFE).
Unfortunately, predictions for the cosmic evolution of metallicity gradients from cosmological simulations and models are limited to 3ā4 so far. If we assumed that the trend with redshift follows a simple linear extrapolation to z 8 of the negative trends from TNG50 (Hemler etĀ al., 2021) and MUGS (Gibson etĀ al., 2013), as well as the model by Kubryk etĀ al. (2015), we would get metallicity gradients that are incompatible with our estimates considering the uncertainties (Fig. 7), which would instead be more in line with flat gradient predictions at high (MaGICC, Gibson etĀ al. 2013; FIRE, Ma etĀ al. 2017; EAGLE, Tissera etĀ al. 2022; Mott etĀ al. 2013 with variable SFE; Sharda etĀ al. 2021). However, this very rough linear extrapolation is most likely wrong, given that metals are expected to start forming from uniformly distributed, pristine gas and therefore the gradients should converge to zero at some point in the past. The only case which extends beyond 3ā4 is the simulations of Taylor & Kobayashi (2017), which reach up to 6, where gradients even steeper than ā1.5 dexĀ kpc-1 (not shown in the figure to avoid an excessive shrink of the axis) are predicted, way steeper than those observed at z 6ā8; however, even in this case, the redshifts up to 8 probed by the current observations are not explored.
All in all, we cannot draw any clear conclusion on what cosmological simulations and models best reproduce the observed gradients at high . Instead, we stress the need for updated metallicity gradient predictions from cosmological simulations and chemical models at 3 in order to match the redshifts reached by current observations.
@@
In summary, the flat radial gradients and the asymmetric spatial distribution of metallicity observed in some cases seem to support the merging scenario, though ISM mixing due to past SN winds or re-accretion in the outer regions of metal-enriched gas from past galactic outflows (galactic fountains) are also a viable mechanism to explain the flatness of the gradients. Spatially asymmetric low-metallicity accretion from the circumgalactic medium (CGM) and intergalactic medium (IGM) could also contribute to produce non-azimuthally symmetric metallicities as suggested by Arribas etĀ al. (2023).
@@ 9 Conclusions
The targets have very disturbed morphologies in both the stellar continuum and ionised gas line emission. We identify several spatial components concentrated within a few kiloparsecs in all the three systems. We obtained the main integrated properties of each spatial component in each target, specifically stellar mass from SED fitting, SFR from H or H, and gas-phase metallicity by means of flux ratios of rest-frame optical emission lines which we modelled with Gaussian functions. We found log() 7.6ā9.3, SFRs 1ā15 Ā yr-1, and gas-phase metallicities 12+log(O/H) 7.7ā8.3 by extracting integrated spectra from circular apertures centred on each spatial component.
We compared the stellar masses and gas-phase metallicities of the targets with the mass-metallicity relations (MZRs) inferred observationally at different redshifts and with those predicted by cosmological simulations. In general, the sources in the systems studied in this work are consistent with the distribution of the highest- galaxies to date ( 3ā10) from JADES, CEERS, and EROs samples in the mass-metallicity plane. Nevertheless, most of the sources in A2744-YD4 and COSMOS24108 lie at the upper end of this 3ā10 distribution, in terms of metallicity at a given stellar mass, and are closer to the best-fit MZRs measured at 2ā3 (e.g. MOSDEF and GLASS-JWST surveys) than to the MZR at 3ā10. Relative to the MZRs from cosmological simulations, the sources in the three systems studied in this work are in good agreement with the slopes predicted at 6ā8 by most simulations.
@@
Cosmological simulations and chemical evolution models make very different predictions regarding the cosmic evolution of metallicity gradients. Some of them predict steeper negative gradients with increasing redshift (e.g. MUGS and TNG50), while others predict median flatter gradients at higher (e.g. MaGICC, FIRE, and EAGLE). Unfortunately, these predictions are generally limited to 3, therefore no conclusions can be drawn on what simulations best reproduce the observed mainly flat gradients at 6ā8 found in this work.
All in all, the results of this work in terms of the MZR and gas-phase metallicity gradients at 6ā8 provide important constraints to guide future cosmological simulations and models. In particular, they urge for specific predictions on the cosmic evolution of metallicity gradients and metallicity maps out to redshift of 8, given that most of such predictions are currently limited to 3.
@@ Acknowledgements.
- Garnett etĀ al. (1997) Garnett, D.Ā R., Shields, G.Ā A., Skillman, E.Ā D., Sagan, S.Ā P., & Dufour, R.Ā J. 1997, ApJ, 489, 63
@@ Appendix A Additional figures for A2744-YD4
[…]
figureNIRSpec PRISM/CLEAR spectra extracted from a circular aperture with radius of 0.15ā³Ā centred at the location of the targets YD6, YD6-[Oāiii]-E, YD4-[Oāiii]-W, and ZD1.
[…]
![[Uncaptioned image]](x25.png)
![[Uncaptioned image]](x26.png)
![[Uncaptioned image]](x27.png)
![[Uncaptioned image]](x28.png)
[…]
figureAdditional maps for A2744-YD4. Highest-resolution continuum map of A2744-YD4 from the original unsmoothed data cube in observed spectral range 1.2-2 m, where the spatial resolution is the highest (top-left). All the other maps are obtained from the (wavelength-dependent) spatially smoothed data cube, as in Fig. 2 (details in Sect. 2.2). First row: H (centre) and [Oāii] (right) integrated fluxes. Second row: [Oāiii]/H (left) and [Oāiii]/[Oāii] (right) line flux ratios. Contours mark the 2ā3 m (observed-frame) continuum.
@@ Appendix B Additional figures for BDF-3299
[…]
figureNIRSpec PRISM/CLEAR spectra extracted from a circular aperture with radius of 0.1ā³Ā centred at the location of the targets BDF-3299-b and BDF-3299-c.
![[Uncaptioned image]](x30.png)
![[Uncaptioned image]](x31.png)
[…]
figureAdditional maps for BDF-3299. As in Fig. 3, the flux maps are obtained from the original data cube, while the line ratio maps from the spatially smoothed data cube (details in Sect. 2.2). Top row: maps of H integrated flux (left), [Oāiii]/H (centre) and [Oāiii]/[Oāii] (right) line flux ratios.
@@ Appendix C Additional figures for COSMOS24108
[…]
figureNIRSpec PRISM/CLEAR spectra extracted from a circular aperture with radius of 0.15ā³Ā centred at the location of the targets COSMOS24108-b, COSMOS24108-c, COSMOS24108-[OIII]-Ea, and COSMOS24108-[OIII]-Eb.
![[Uncaptioned image]](x33.png)
![[Uncaptioned image]](x34.png)
![[Uncaptioned image]](x35.png)
![[Uncaptioned image]](x36.png)
![[Uncaptioned image]](x37.png)
[…]
figureAdditional maps for COSMOS24108. As in Fig. 4, the flux maps are obtained from the original data cube, while the line ratio maps from the spatially smoothed data cube (details in Sect. 2.2). Top row: median continuum map in the observed spectral range 3ā3.5 m (0.39ā0.46 m rest-frame; left), with the contours of the continuum in the observed spectral range 1ā2 m (same as in Fig. 4, left) superimposed; H (centre-left), [Oāii] (centre-right), and H (right) flux maps. Middle row: [Oāiii]/H (left) and [Oāiii]/[Oāii] (centre-left) line ratio maps.
![[Uncaptioned image]](x39.png)
![[Uncaptioned image]](x40.png)
![[Uncaptioned image]](x41.png)
![[Uncaptioned image]](x42.png)
![[Uncaptioned image]](x43.png)
Metallicity gradients for the two minor sources in COSMOS24108, [Oāiii]-Ea and [Oāiii]-Eb (top and mid panels, respectively), and for the main system (bottom panel), when centring the gradient around the peak of the [Oāiii] emission instead of around the continuum peak as in Fig.Ā 6.